### v1

# SubGrapher: Visual Fingerprinting of Chemical Structures

[…]
SubGrapher: Visual Fingerprinting of Chemical Structures
1 Author list
Lucas Morin1, 2, *, Gerhard Ingmar Meijer1, Valéry Weber1, Luc Van Gool3, 2, Peter W. J. Staar1
2 Affiliations
1 IBM Research, Säumerstrasse 4, 8803 Rüschlikon, Switzerland
2 Department of Information Technology and Electrical Engineering, ETH Zürich, Sternwartstrasse 7, 8092 Zürich, Switzerland
3 INSAIT, Sofia University St. Kliment Ohridski, Tsarigradsko shose  [^1] R, 1784 Sofia, Bulgaria
* Corresponding author:
Lucas Morin, lum@zurich.ibm.com
3 Keywords
Molecular Fingerprint, Optical Chemical Structure Recognition, Document Understanding, Functional Group Recognition
4 Abstract
Automatic extraction of chemical structures from scientific literature plays a crucial role in accelerating research across fields ranging from drug discovery to materials science. Patent documents, in particular, contain molecular information in visual form, which is often inaccessible through traditional text-based searches. In this work, we introduce SubGrapher, a method for the visual fingerprinting of chemical structure images. Unlike conventional Optical Chemical Structure Recognition (OCSR) models that attempt to reconstruct full molecular graphs, SubGrapher focuses on extracting molecular fingerprints directly from chemical structure images. Using learning-based instance segmentation, SubGrapher identifies functional groups and carbon backbones, constructing a substructure-based fingerprint that enables chemical structure retrieval. Our approach is evaluated against state-of-the-art OCSR and fingerprinting methods, demonstrating superior retrieval performance and robustness across diverse molecular depictions. The dataset, models, and code will be made publicly available.
5 Introduction
Knowledge in chemistry is spread across structured databases and unstructured sources such as scientific journals, patent documents, books, and corporate documents. Integrating molecular structures, properties, and synthesis data into a unified collection would significantly accelerate research in chemistry and materials science [1]. Converting unstructured documents to machine readable format enables searching within document collections and training foundational models at larger scales [2, 3]. A key challenge in this process is extracting molecular structures from patent documents, where they are primarily represented as images rather than machine-readable text. Molecular databases such as PatCID [4] and SureChEMBL [5] aim to address this challenge by employing document ingestion pipelines that integrate page segmentation, chemical image classification, and Optical Chemical Structure Recognition (OCSR).
The main component of these pipelines, OCSR, is the process of converting molecular depictions into structured representations, such as SMILES (Simplified Molecular Input Line Entry System) [6] or molecular graphs. Graph-based OCSR methods identify molecular components, including atoms and bonds, and reconstruct their connectivity. These methods rely on rule-based image processing algorithms, as seen in OSRA [7], Imago [8], and MolVec [9], or deep-learning models such as MolGrapher [10] and ChemGrapher [11]. In contrast, sequence-based OCSR methods, such as DECIMER [12] and MolScribe [13], utilize vision encoders with autoregressive text decoders to generate SMILES strings.
Despite advances, OCSR methods face challenges with variations in chemical drawing conventions or degraded image quality. Additionally, certain chemical illustrations cannot be represented as SMILES, particularly those containing non-standard graphical elements (see row 3 in Figure 4) or Markush structures [14]. This limitation is especially significant in patent analysis, where Markush structures are widely used to define broad molecular classes. Moreover, full molecular reconstruction is not always necessary for applications such as database searching or molecular property prediction. In many cases, users are more interested in identifying molecules with specific substructures rather than their complete structures. Furthermore, predictive models often rely on molecular fingerprints derived from SMILES [15], making SMILES an intermediate representation rather than a final output.
Figure 1: SubGrapher extracts a molecular fingerprint from a chemical structure image in a document. Our approach identifies functional groups and carbon backbones in chemical structure images. These substructures are then combined based on their connectivity to create a substructure-graph. Finally, this graph is converted to a fingerprint enabling substructure search, similarity search or any downstream predictive task.
Molecular fingerprints are vectorized representations of molecular structures, commonly used for similarity searches in databases and predictive tasks such as property prediction or drug target identification [16]. Substructure-based fingerprints, such as MACCS [17], PubChem Fingerprints [18], RDKit Daylight [19], and Extended-Connectivity Fingerprints (ECFP) [20], encode the presence or absence of predefined molecular fragments. Among these substructures, functional groups [21] form a key subset due to their well-defined chemical properties. Besides, molecular fingerprints can also be generated using probabilistic methods like MHFP [22] or learned from large datasets using models such as MoLFormer [23].
Integrating OCSR and fingerprinting into a single-step process by directly recognizing functional groups from images would efficiently facilitate the extraction of molecular information from literature. Recently, Fan, et al., [24] explored the recognition of a limited subset of functional groups from images as an auxiliary task to enhance OCSR performance. Additionally, other studies [25, 26] have investigated the use of molecule images to learn molecular features used for predictive tasks. However, none of these studies have combined visual recognition of functional groups with visual fingerprinting into a unified approach.
In this work, we introduce SubGrapher for the visual fingerprinting of chemical structure images, illustrated in Figure 1. First, SubGrapher uses segmentation models to identify functional groups and carbon backbones in images. Second, SubGrapher creates a substructure-graph based on the connectivity of these substructures. Finally, this graph is converted to a molecular fingerprint. On the one hand, our fingerprint allows to search for molecules containing functional groups and carbon backbones of interest. By defining key chemical properties, functional groups are natural targets for molecular searches, while carbon backbones link these functional groups. On the other hand, SubGrapher’s fingerprint can be used for approximate search of exact chemical structures.
6 Method
We introduce SubGrapher, a model designed for the visual fingerprinting of chemical structure images. Its architecture is illustrated in Figure 2. SubGrapher employs substructure segmentation models to detect and identify molecular substructures, specifically functional groups and carbon backbones. These detected substructures are then assembled into a substructure-graph, which is subsequently converted into a molecular fingerprint.
Figure 2: SubGrapher architecture. The SubGrapher model detects functional groups and carbon backbones using instance segmentation networks. The identified substructures are combined into a substructure-graph, which is then converted to a matrix fingerprint.
6.1 Substructure Segmentation Model
To detect molecular substructures from images, SubGrapher employs two segmentation networks, as illustrated in Figure 2. The first network detects 1,534 expert-defined functional groups, ensuring broad coverage of molecular diversity. The second network identifies 27 distinct carbon backbone patterns, complementing the functional group detection by capturing molecular regions not already assigned to any functional groups. By ensuring that substructures overlap appropriately and span the entire molecule, SubGrapher constructs a more characteristic and informative fingerprint in later stages.
Substructures are detected using mask-based segmentation rather than relying solely on bounding-box detection. This approach provides fine-grained supervision during training, leading to improved accuracy and robustness in molecular substructure recognition. Additionally, our framework enables training substructure recognition directly from page images, eliminating the need for segmentation in document-to-chemical-structure pipelines [4]. In the implementation, the segmentation models are Mask-RCNN [27] models.
6.2 Substructure-Graph Construction
The detected substructures are used to construct a substructure-molecular graph, where nodes represent the identified substructures and edges correspond to their intersections.
Substructures are considered intersected if their predicted bounding boxes overlap. To enhance connectivity, each detected box is expanded by a fixed amount. This expansion ensures that adjacent substructures sharing a single atom are linked. Additionally, substructures that are not directly connected in the molecular graph may still form connections in the substructure-graph if they are spatially close in the depicted molecular conformation. This added connectivity incorporates positional information, improving the representativity of the substructure graph. In the implementation, the expansion of each bounding box is set to 10% of the diagonal length of the smallest detected box in the image. This constraint prevents excessive expansion, ensuring a reasonable balance that avoids artificially distant intersections.
6.3 Substructure-based Visual Molecular Fingerprint
To describe molecular structures, SubGrapher constructs a visual molecular fingerprint by combining detected substructures. This fingerprint encodes the occurrences of functional groups and carbon segments, as well as their distances.
We introduce the substructure-based visual molecular fingerprint (SVMF). For a given molecule $m$, SVMF$(m)$ is defined as an upper triangular matrix:
$\text{SVMF}(m)=\begin{bmatrix}&\vspace{-2mm}\ddots&&&&&&&\\ &&&\hskip 5.69054ptf_{i,i}&\hskip 5.69054pt\dots&\hskip 5.69054ptf_{i,j}&&&\\ &&&&\vspace{-2mm}\ddots&\hskip 5.69054pt\vdots&&&\\ &&&{\text{0}}&&\hskip 5.69054ptf_{j,j}&&&\\ &&&&&&\ddots&&\\ \end{bmatrix}$ (1)
[…]
$f_{i,i}=h_{1}.n_{i}$ (2)
where $n_{i}$ is the number of instances of substructure $i$ in the molecule, and $h_{1}$ is a model hyperparameter. The off-diagonal elements, $f_{i,j}$, represent the intersection coefficient between substructures $i$ and $j$:
$f_{i,j}=\sum_{\alpha=1}^{n_{i}}\sum_{\beta=1}^{n_{j}}h_{2}\circ d(i_{\alpha},j_{\beta})$ (3)
where $\circ$ is the composition operator, $d(i_{\alpha},j_{\beta})$ denotes the distance between instances $i_{\alpha}$ and $j_{\beta}$, and $n_{i}$ and $n_{j}$ are the number of instances of the substructures $i$ and $j$. The distance $d(i_{\alpha},j_{\beta})$ is defined as the smallest number of substructures required to construct a path between instances $i_{\alpha}$ and $j_{\beta}$. The function $h_{2}$ is an hyper-parameter of the model controlling the number of non-zero intersection coefficients, and ultimately the fingerprint compression.
In the implementation, $h_{1}$ is set to 10, while $h_{2}(d)$ takes values of $2$, $2/4$, $2/16$, and $2/256$ for distances $d=1,2,3,4$, respectively. If $d>4$, the intersection coefficient is set to zero. For carbon chain instances, intersection coefficients are divided by $2$ to assign greater significance to functional groups. This prioritization reflects the intuition that functional groups are the most important parts of the molecule. The SVMF matrix can be transformed into a one-dimensional fingerprint by first converting all values to integers using a fixed precision cutoff and then separately encoding active matrix positions and their corresponding values.
6.4 Molecule Retrieval
The similarity between two fingerprints of molecules $m_{1}$ and $m_{2}$, $f(m_{1})$ and $f(m_{2})$, can be computed as:
$s(f(m_{1}),f(m_{2}))=\frac{\parallel f(m_{1})-f(m_{2})\parallel}{\parallel f(m_{1})+f(m_{2})\parallel}$ (4)
where $\parallel$.$\parallel$ . ∥ denotes the Euclidian norm. This similarity measure enables molecular searches in large datasets. Given a query molecule, users can retrieve similar structures. Further details are in the Results section below.
6.5 Substructures Coverage
Figure 3: Substructure examples. Examples of functional groups recognized by SubGrapher.
Here, we describe the strategy used to select the 1,534 functional groups and 27 carbon backbone patterns.
Our functional groups are defined as substructures containing at least one heteroatom and having for attachment points single bonds connected to carbon atoms. These substructures are manually designed starting from chemically logical chains of up to five atoms selected from C, O, S, N, B, or P. Each heteroatom in these initial chains can be expanded with chemically relevant subgroups containing atoms C, H, O, S, or N. For instance, if sulfur is present in the primary chain, we also consider replacing it with SO and SO2. To refine the selection, we perform substructure searches in PubChem for candidate ‘functional group family’ (i.e., chemically similar groups). Families with fewer than approximately 1,000 occurrences in PubChem were not further expanded. Some example functional groups are illustrated in Figure 3.
For carbon backbones, we include standard combinations of three to six carbon atoms with single, double, or triple bonds, as well as common three- to six-membered rings. If a carbon backbone substructure is fully contained within another substructure, we retain only the larger one to minimize redundancy. We defined functional groups and carbon backbones substructures to ensure that they are sufficiently overlapping. This allows our molecular fingerprint to remain distinctive and applicable to a wide range of organic compound.
6.6 Training Dataset Generation
Datasets of molecular images extracted from real documents are scarce. Especially, there is no dataset which includes mask annotations. To address this limitation, we use the synthetic data generation pipeline proposed in MolDepictor [10], which generates a diverse set of molecular depictions. This pipeline selects molecular SMILES from PubChem and renders them using the RDKit library [19].
As this pipeline was originally designed to generate molecular graphs for supervision, we extend it to produce mask annotations for detected substructures. To generate an image for each substructure within a molecule, we first identify substructures using RDKit. Next, we post-process the generated SVG images, mapping SVG elements to the corresponding atoms and bonds within each substructure.
Additionally, we improve the pipeline to generate depictions of Markush structures, including structural, positional, and frequency variation indicators. For this purpose, we first construct artificial Markush structures represented as CXSMILES [28], modifying SMILES from PubChem by replacing atom labels with R-group labels, attaching functional groups or R-groups to rings, and including bracket notations. Finally, we use the CDK library [29] to generate Markush structure depictions along with their corresponding substructure mask annotations.
7 Results and Discussion
In this section, we perform experiments to evaluate SubGrapher for substructure detection and molecular fingerprinting. Our method is evaluated against state-of-the-art OCSR models and SMILES-based fingerprinting methods. Moreover, we analyze the strengths and weaknesses of SubGrapher on qualitative examples.
7.1 Substructure Detection Evaluation
7.1.1 Datasets and Metrics
To compare our method with state-of-the-art approaches, we evaluate it on three benchmarks: JPO [30], a subset of USPTO-10K-L [10], and USPTO-Markush [31]. These datasets contain pairs of chemical images and molecular structures stored as MOL files [32].
JPO consists of 341 molecule images extracted from patent documents published by the Japanese Patent Office [33]. This dataset is challenging due to non-standard drawing conventions and low image quality. Molecules containing abbreviations were removed from the standard JPO benchmark set. USPTO-10K-L contains 10,000 images of large molecules with over 70 atoms and no abbreviations. We evaluate our method using the first 1,000 images of this benchmark. USPTO-Markush includes 74 images of Markush structures from patent documents published by the United States Patent and Trademark Office (USPTO [34]).
MOL files [32] contain the atoms, bonds and their connectivity, allowing us to extract the ground-truth substructures present in a molecule. The substructures evaluated are the 1534 functional groups and the carbon backbones are not considered.
To evaluate performances, we compute the Substructure F1-score (S-F1) and the Molecule Exact Match (M-EM). The substructure F1-score measures recall (the proportion of ground-truth substructures correctly identified by predictions) and precision (the proportion of predicted substructures that correctly match the ground truth) for substructure detection. The molecule exact match is the percentage of molecules where all substructures are perfectly detected, i.e., the S-F1 equals one. Since OCSR models such as OSRA and DECIMER do not predict atoms positions, our evaluation metrics consider only the presence or absence of substructures, not their exact location.
7.1.2 State-of-the-art Comparison
Methods JPO (341) USPTO-10K-L (1000) USPTO-Markush (74)
S-F1 M-EM       S-F1 M-EM       S-F1 M-EM
Rule-based methods
OSRA [7] 81 67       97 75       74 70
Learning-based methods
DECIMER [35] 86 79       86 66       10 11
MolScribe [36] 94 82       90 55       86 86
MolGrapher [10] 89 80       56 31       35 30
SubGrapher (Ours) 92 83       97 55       88 82
Table 1: Substructures detection comparison. Evaluation on benchmarks of molecule images (JPO, USPTO-10K-L) and Markush structure images (USPTO-Markush). Substructure F1-score (S-F1) evaluates the recall and precision for substructure detection. Molecule Exact Match (M-EM) evaluates the percentage of molecules where all substructures are perfectly detected, i.e., the S-F1 equals one. Best scores are bold and second-best scores are underlined.
Figure 4: Qualitative evaluation. Examples of predicted functional groups are shown for images from patent documents (JPO, USPTO-10K-L, USPTO-Markush) and a scientific journal (JOC).
Table 1 compares SubGrapher with state-of-the-art methods for substructure detection. Our method achieves the highest molecule exact match on the JPO dataset. It demonstrates its robustness to the lower-quality images commonly found in this dataset.
On the USPTO-10-L dataset, SubGrapher outperforms other deep-learning approaches in substructure F1-score. Unlike models such as MolGrapher and MolScribe, which experience a performance drop with larger molecules, SubGrapher maintains consistent performance. This suggests that, compared to OCSR models, object detectors are better equipped to handle variations in image scale. Our method is also supervised using a stronger supervision than standard OCSR models, specifically pixel-level (mask) annotations rather than supervision from SMILES strings or molecular graphs. This enables improved recognition of fine details in images of large molecules.
On the USPTO-Markush dataset, SubGrapher again outperforms other methods in substructure F1-score. This superior performance results from the limitations of competing models: OSRA and MolGrapher are trained exclusively on molecular images, whereas MolScribe and DECIMER handle Markush structures but support only limited Markush features, specifically variable groups represented as abbreviations. In contrast, SubGrapher effectively manages complex Markush structures by focusing on relevant regions of the image and excluding Markush-related annotations.
7.2 Qualitative Evaluation
Figure 4 showcases examples of predicted molecules for images from various benchmark datasets. SubGrapher accurately recognizes functional groups in molecule images that contain captions or are of lower quality (Figure 4, row 1). It also recognizes functional groups in complex Markush structures (Figure 4, row 2) or unconventional drawings displayed in scientific publications (Figure 4, row 3). Additionally, our method maintain strong performances on extremely large molecules (Figure 4, row 4). Unlike image captioning methods such as DECIMER, our predictions preserve the spatial arrangement of substructures from the input image. This conveys useful information for human interpretation.
7.3 Visual Fingerprinting Evaluation
In this section, we evaluate SubGrapher and state-of-the-art methods for the retrieval of a molecule in a collection of molecule images.
Figure 5: Visual fingerprinting evaluation strategy. (A) First, a set of similar molecule SMILES is rendered into molecular images. These images are subsequently converted into fingerprints using each of the evaluated methods. (B) Second, a query SMILES string is transformed into a fingerprint. Its similarity is computed against all fingerprints within the dataset. Finally, based on the ranking of the similarity scores, the position of the correct molecule is reported.
7.3.1 Datasets and Metrics
We evaluate visual fingerprinting methods using five benchmark datasets: adenosine, camphor, cholesterol, limonene, and pyridine. The construction of these benchmarks is illustrated in Figure 5.
First, for each reference molecule (adenosine, camphor, cholesterol, limonene, and pyridine), we sample 500 SMILES from PubChem that have at least 90% structural similarity to the corresponding reference molecule. Second, we convert this sets of similar molecules to a sets of images by rendering them using RDKit. The resulting images undergo substantial data augmentation, including shifting, scaling, rotation, downscaling, and grid distortion. Third, each visual fingerprinting method converts these augmented images into molecular fingerprints. The methods evaluated are SubGrapher, OSRA with RDKit Daylight [19] or MHFP [22], and MolScribe with RDKit Daylight or MHFP. Then, to assess retrieval performance, we measure each method’s ability to correctly retrieve a molecular image using its fingerprint from a provided SMILES query. First, the query SMILES is converted into a fingerprint based on the evaluated method. Second, the similarity between this query fingerprint and all fingerprints in the dataset is computed. Third, the results are ranked based on similarity scores, and the position of the correct ground-truth molecule is recorded. Each benchmark is tested using 50 SMILES queries, and the average rank across these queries is reported. Intuitively, this setup evaluates the ability to retrieve a molecular image from a document based on a user-provided SMILES query under extreme conditions, as all molecular images in the dataset are highly similar and augmented. Although the test sets contain only 500 images, this design aims to simulate searches in larger collections.
7.3.2 State-of-the-art Comparison
Models Adenosine (500) Camphor (500) Cholesterol (500) Limonene (500) Pyridine (500)
OCSR Fingerprint
OSRA [7] RDKit [19] 184 148 165 219 210
MolScribe [13] RDKit 217 363 149 272 203
OSRA MHFP [22] 139 104 181 152 114
MolScribe MHFP 101 287 187 48 283
SubGrapher (Ours) 110 73 106 81 103
Table 2: Visual fingerprinting comparison. We compare the retrieval performance of various Optical Chemical Structure Recognition (OCSR) and fingerprinting methods on image datasets generated from adenosine, camphor, cholesterol, limonene, and pyridine. We report the average rank at which a query molecule is retrieved across 50 queries. Best scores are bold and second-best scores are underlined.
Table 2 compares visual fingerprinting methods for similarity search. On average, SubGrapher ranks first for camphor, cholesterol, and pyridine, and second for adenosine and limonene. Additionally, SubGrapher retrieves the correct reference molecule at an average rank of 95, significantly outperforming other methods. Given the challenging nature of the evaluation, we expect that, in practical scenarios, the target molecule would typically appear within the top 100 results, making manual inspection feasible. SubGrapher also performs more consistently across different types of reference molecules, suggesting that it handles structural variations reliably. Our results show that converting molecular images directly into fingerprints in a single step produces more distinctive fingerprints.
8 Conclusion
We present SubGrapher a learning-based approach for converting 2D chemical structure depictions to molecular fingerprints. SubGrapher detects functional groups and carbon backbones, assembling them into a graph representation, which is subsequently converted into a fingerprint. This fingerprint enables efficient substructure and molecule retrieval from collections of molecular images. Trained on synthetic images, SubGrapher demonstrates strong generalization capabilities, making it useful for large-scale molecular search. Our results highlight the benefits of strong mask supervision in molecular image recognition and demonstrate that directly converting images to fingerprints in a single step significantly improves molecular retrieval performance.